two liquid A and B at temperature 75 celcius and 15 celsius their masses are in ratio 2:3 and specific heat is in ratio 3:4 .what is their resulting temperature?

Grade:12

2 Answers

Adarsh
6 years ago
Dear Isha ,
The final temperature of mixture will be 350C.
Thanks and Regards 
Adarsh
Khimraj
6 years ago
Let mass of Liquid A and Liquid B is ma and mb
specific heat of Liquid A and Liquid B is sa and sb
and Final temperature be T0 C Then
ma/mb = 2/3 and sa/sb = 3/4
Qloss = Qgain
masa(75-T) = mbsb(T-15)
(ma/mb)(sa/sb) = (T-15)/(75-T)
(2/3)(3/4) = (T-15)/(75-T)
½ = (T-15)/(75-T)
75-T = 2T – 30
3T = 105
T = 350C.
Hope it clears.